OpenStudy (anonymous):

what's 2y^2-9y+20.25 factored out?

OpenStudy (jhannybean):

I would suggest using the completing the square method.

OpenStudy (anonymous):

yeah im using that method rn

OpenStudy (anonymous):

the original equation was 2y^2=9y-9

OpenStudy (jhannybean):

then how did you get 20.25 from that?

OpenStudy (jhannybean):

all you have to do there is move the \(9y-9\) over to the left.

OpenStudy (jhannybean):

So it would look like \[2y^2-9y-9=0\] instead of \[2y^2 -9y +20.25\]

OpenStudy (anonymous):

i moved 9y to the left and i got 2y^2-9y=-9

OpenStudy (jhannybean):

Okay and now move the \(\-9\) over to the left as well.

OpenStudy (anonymous):

then i took half of the b coefficient, squared it and added it to both sides

OpenStudy (jhannybean):

Did you also factor out the 2 like this? : \[2\left(y^2-\frac{9}{2}y\right)-9=0\]

OpenStudy (jhannybean):

When you factor out the 2, it makes it easier to create a quadratic inside the parenthesis. You are essentially completing the square inside

OpenStudy (jhannybean):

Are you understanding what I did?

OpenStudy (anonymous):

uhm totally

OpenStudy (jhannybean):

Where are you confused first? Let's clarify that.